OpenStudy (anonymous):

What are the solutions of 4x2 + x = -3?

OpenStudy (anonymous):

the quantity negative 1 plus or minus 7i all over 8 the quantity negative 1 plus or minus i square root of 5 all over 2 the quantity negative 1 plus or minus i square root of 47 all over 8 the quantity negative 1 plus or minus 2i square root of 7 all over 8

OpenStudy (anonymous):

4x2 + x +3=0

OpenStudy (anonymous):

now use quadratic formula

OpenStudy (anonymous):

for ax^2+bx+c =0 x = (-b + - sqrt(b*b -4*a*c) )/ ( 2*a) here a=4,b=1,c=3

OpenStudy (anonymous):

okay, so what I got was 

OpenStudy (anonymous):

@matricked